OpenStudy (anonymous):

is sin(x^2) the same as sin^2x

OpenStudy (allank):

I believe so.

OpenStudy (anonymous):

just to make it clearer \[\sin(x^2)=\sin^2x?\]

OpenStudy (allank):

Still believe that's true.

OpenStudy (australopithecus):

no \[\sin^2(x) = \sin(x)\sin(x)\] \[\sin(x^{2}) = \sin(x^{2}) \]

OpenStudy (australopithecus):

how about we sub in a value to see if it is true

OpenStudy (anonymous):

so how would i integrate \[\sin(x^2)\]

OpenStudy (australopithecus):

put in pi/4 into both functions and see if they give you the same output

OpenStudy (australopithecus):

as I suspected https://www.wolframalpha.com/input/?i=sin%28%28pi%2F4%29^2%29+%3D+sin^2%28pi%2F4%29